OpenStudy (anonymous):

Probability Question: Ten boys and 12 girls decide to rent a 16-passenger van and a six-passenger car to drive to a rock concert in a nearby city. If the group is distributed randomly between the vehicles, what is the probability that a) there are no boys in the car? b) there are no girls in the car? c) Alan and Margaret are both in the van? d) there are more girls than boys in the car? I just need help with C and D please...

OpenStudy (kamille):

So,have you heard about using C?

OpenStudy (anonymous):

yes.

OpenStudy (kamille):

ah, I am lost, let's invite smarter people @amistre64

OpenStudy (kamille):

@hartnn

OpenStudy (amistre64):

C is not discernable without an assumption

OpenStudy (amistre64):

how did you determine a and b? the process is similar for c and d

OpenStudy (anonymous):

Here's what I did for letter C: Since Margaret and Alan is already in the van, then there will only be 20 people left. And the available sits in the van will be 20. So event A= 20C14 (20 sits for 14 people). Then \(P(A)=\frac{20 C 14}{22C16}\) then I got the right answer. but i'm not sure if my reasoning makes sense.. For letter A: 12C6/22C6 For letter B: 10C6/22C6

OpenStudy (amistre64):

gggbbb is equal so ggggbb or gggggb or gggggg the sum of the individual probabilities gives us P([4,2] or [5,1] or [6,0])

OpenStudy (anonymous):

you're doing letter D?

OpenStudy (amistre64):

that was the concept for D yes

OpenStudy (anonymous):

oh okay (: i'm also doing it. what will be the sample space for this one?

OpenStudy (amistre64):

the sample space for C im still trying to think of :) Zarkon or Satellite prolly can do it in thier sleep

hartnn (hartnn):

since they 2 are already in the van 2 seats are filled only 4 remains

OpenStudy (amistre64):

12P4 * 10P2 out of 22P6 = 4g 2b in the car right? P or C work the same in this case

hartnn (hartnn):

i am bit confused though 6 passenger or 16 passenger ?? you used 16 for C but 6 for A,B ?

OpenStudy (anonymous):

16 passenger seats for the van and 6 for the car

hartnn (hartnn):

ok, so for C only, we consider 16

OpenStudy (anonymous):

P and C are different though, aren't they? O.o @amistre64

OpenStudy (anonymous):

out of 22 sits (from car & van), 16 sits are from the van so 22C16 is what i used for my sample space in letter C

OpenStudy (amistre64):

they are are related: k! nCk = nPk but i think i recall if all the terms are of the same 'kind' that we are fine

hartnn (hartnn):

why 22 again ? those 2 have fixed their seats so 16 choices out of 20 remaining

hartnn (hartnn):

ok, ignore me you're finding sample space

hartnn (hartnn):

so 20C14 / 22 C16 is correct what was the doubt ?

OpenStudy (anonymous):

so it's right? nothing, i just wanted to confirm (: how about sample space in D?

OpenStudy (amistre64):

i think C instead of P is better for my thought on d, tried some stuff with the wolf and failed at it lol

hartnn (hartnn):

amistre's logic is correct for D and yes, it should be nCr only

OpenStudy (anonymous):

the right answer for D is 0.42 but I got 0.042 here's what I did: I use Direct method: (12C6 x 10C0)+(12C5 x 10C1) + (12 C 4 x 10 C 2) =3119 Sample Space: 22C6 (because they are talking about the car) so 3119/22C6=0.042

OpenStudy (amistre64):

31119/(22C6)

OpenStudy (anonymous):

oh right! i'm missing one "1". Thanks ! @amistre64 & @hartnn

OpenStudy (amistre64):

good luck ;)

hartnn (hartnn):

thanks for asking good questions :)